What would the balance look like if you replace the unknown weight with an 8-ounce weight? The left side is 6oz the right side is 2oz with the unknown weight. Also explain why the solution of 2+x=6 cannot be that x is equal to 8
1
Expert's answer
2020-04-23T17:43:09-0400

Left side would be downward, because 2+8=10>6.

2+x=6

x=6-2

x=4

It's obvious that 8 and 4 are not equal.